R. Rz R3 21 20 it 70 Vje R5 94 35 I II Find Vaat and R₃

value of r3 to make Vx=0
Answer #1

R₂ VA 21. 2012 Is R4 Rs I VA 94 35 Iz A 17A find the value R₂ make Vu=0 R R₂ & R₂ are series RG = R + R2 + R₂ = 20+21+ R₂ R₂from circuit A find voltage VA SVE Is (R; + R2) W A 2193 (170 +R3) 21+20) V 89913 170 + R₂ 17 (41+R₂) Iz Ry Xg4 (170+R3 ) Veput equation 1741+70) Iz 7.8625 A 17.0+ 70 from equation VA = I6 (R, +R₂) - 9. 1375 221 +20) = 37.4.637.5 Volt Iz Ry = 718625